# Learning to Explore with Lagrangians for Bandits under Unknown Linear Constraints

### v1

@@ H.1 Upper Bound on Constraint Violation
In a linear programming problem we say constraint is violated if the chosen allocation fails to satisfy any of the true linear constraints. In other words when the event $\mathbf{A}\boldsymbol{\omega}_{t}\geq 0$. We start with the optimization problem relaxed with slack if the constraints were known,
$\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}{\boldsymbol{\mu}}^{T}\boldsymbol{\pi}$
such that, $\mathbf{A}\boldsymbol{\pi}+\Gamma\leq 0$
where, $\Gamma$ is the slack. Cumulative violation of constraints can be expressed as $\mathcal{V}_{t}=\sum_{s=1}^{t}\max_{i\in[K]}\left[\mathbf{A}^{i}\boldsymbol{\omega}_{t}\right]_{+}$ where $[z]_{+}=\max\{z,0\}$. Then, at any time step $t\in[T]$, instantaneous violation is given by, $v_{t}=\max_{i\in[k]}\left[\mathbf{A}^{i}\boldsymbol{\omega}_{t}\right]_{+}$. Since, $\mathbf{A}$ is feasible, we define the game value as, $\eta=\max_{\boldsymbol{\omega}\in\mathcal{F}}\max_{i\in[K]}\mathbf{A}^{i}\boldsymbol{\omega}\leq-\Gamma$ and $\eta=\max_{\boldsymbol{\omega}\in\mathcal{F}}\max_{i\in[K]}\mathbf{A}^{i}\boldsymbol{\omega}\geq\min_{\tilde{\mathbf{A}}\in\mathcal{C}_{A}^{t}}\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}_{t}}\max_{i\in[K]}\tilde{\mathbf{A}}_{t}^{i}\boldsymbol{\omega}_{t}=\tilde{\mathbf{A}}_{t}^{i_{t}}\boldsymbol{\omega}_{t}$, holds because of pessimistic estimate of $\mathbf{A}$.
Again, we define, $i_{\text{min}}(\boldsymbol{\omega})=\arg\min_{i\in[K]}\tilde{\mathbf{A}}^{i}\boldsymbol{\omega}$.
$\max_{i\in[K]}\left[\mathbf{A}^{i}\boldsymbol{\omega}_{t}\right]$ $=\max_{i\in[K]}(\mathbf{A}^{i}-\tilde{\mathbf{A}}_{t}^{i})\boldsymbol{\omega}_{t}+\max_{i\in[K]}\tilde{\mathbf{A}}_{t}^{i}\boldsymbol{\omega}_{t}$
$\leq\max_{i\in[K]}||(\mathbf{A}^{i}-\tilde{\mathbf{A}}_{t}^{i})||_{\Sigma_{t}}||\boldsymbol{\omega}_{t}||_{\Sigma_{t}^{-1}}+\max_{i\in[K]}\tilde{\mathbf{A}}_{t}^{i}\boldsymbol{\omega}_{t}$
$\leq f(t,\delta)||\boldsymbol{\omega}_{t}||_{\Sigma_{t}^{-1}}+\max_{i\in[K]}\tilde{\mathbf{A}}_{t}^{i}\boldsymbol{\omega}_{t}$
$\leq\rho(t,\delta)-\Gamma$
where the last inequality holds because $\max_{i\in[K]}\tilde{\mathbf{A}}_{t}^{i}\boldsymbol{\omega}_{t}\leq\max_{i\in[K]}\mathbf{A}^{i}\boldsymbol{\omega}_{t}\leq-\Gamma$
@@
$\mathbf{E}\big{[}\frac{\mathcal{V}_{\tau_{\delta}}}{\tau_{\delta}}\big{]}=\sum_{t\leq\mathbf{E}[\tau_{\delta}]}s_{t}\leq$ $\sum_{t\leq\mathbf{E}[\tau_{\delta}]}[\rho(t,\delta)-\Gamma]_{+}$
$=$ $\sum_{t\leq\mathbf{E}[\tau_{\delta}]}\{\rho(t,\delta)-\Gamma\}\hskip 5.69046pt\text{because, }\rho(t,\delta)-\Gamma\hskip 2.84544pt\text{ ispositive.}$.
@@ H.2 Experimental results on constraint violation
We compare the constraint violation i.e the number of times $\mathbf{A}\boldsymbol{\omega}_{t}>0$, where $\mathbf{A}$ is the true constraint matrix.
Figure 6: Constraint violation (median±$\pm$±std.) algorithms for hard environment in setup 1.
Figure 7: Constraint violation (median±$\pm$±std.) of algorithms for easy environment in setup 1.
Figure 8: Constraint violation (median±$\pm$±std.) algorithms for hard environment in setup 2.
Figure 9: Constraint violation (median±$\pm$±std.) of algorithms for easy environment in setup 2.
Figure 10: Constraint violation (median±$\pm$±std.) of algorithms for IMDB environment.
Observation: LAGEX is "safest".
@@
Therefore we see that across multiple environments LAGEX has proven to be the "safest" out of all the competing algorithms.
@@ Appendix I Experimental details
Computation resources. We run the algorithms on a 64-bit 13th Gen Intel® Core™ i7-1370P × 20 processor machine with 32GB ram with a disc capability of 1TB and graphics Mesa Intel® Graphics (RPL-P). The hardware model is HP EliteBook 840 14 inch G10 Notebook PC. Code is available in this Link.
@@
Abbreviations of algorithms. "CTnS-WLag" and "CGE-WLag" is basically the CTnS and CGE algorithm (Carlsson et al., (2024)) under unknown constraints without Lagrangian relaxation. "PTnS" is Projected Track and Stop algorithm.
Table 1: Different experimental setups
Setup Environment $\mu$ Constraints $\delta$
$1$ Hard $[1.5,1.0,0.5,0.4,0.3,0.2,0.1]$ $\boldsymbol{\pi}_{1}+\boldsymbol{\pi}_{2}+\boldsymbol{\pi}_{3}\leq 0.5$ and $\boldsymbol{\pi}_{4}+\boldsymbol{\pi}_{5}\leq 0.5$ $0.01$
$1$ Easy $[1.5,1.0,1.3,0.4,0.3,0.2,0.1]$ $\boldsymbol{\pi}_{1}+\boldsymbol{\pi}_{2}+\boldsymbol{\pi}_{3}\leq 0.5$ and $\boldsymbol{\pi}_{4}+\boldsymbol{\pi}_{5}\leq 0.5$ $0.01$
$2$ Hard $[1,0.5,0.4,0.4,0.5]$ $\boldsymbol{\pi}_{1}+\boldsymbol{\pi}_{2}\leq 0.5$ and $\boldsymbol{\pi}_{3}+\boldsymbol{\pi}_{4}\leq 0.5$ $0.1$
$2$ Easy $[1,0.5,0.4,0.95,0.8]$ $\boldsymbol{\pi}_{1}+\boldsymbol{\pi}_{2}\leq 0.5$ and $\boldsymbol{\pi}_{3}+\boldsymbol{\pi}_{4}\leq 0.5$ $0.1$
$\times$ IMDB $[3.67,2.97,2.94,3.52,3.18,2.02,2.79,2.96,2.37,2.53,2.55,2.54]$action" < 0.3, "drama" > 0.3, "family" > 0.3 0.1
I.1 Sample complexity plots for Setup 2
Setup 2. We further test our proposed algorithms LATS and LAGEX under the same two environments used in Figure 3b and 3c of  (Carlsson et al.,, 2024). Figure 12 and 12 shows the comparison of sample complexity (Median $\pm$ s.d) for LATS and LAGEX with other baseline algorithms.
Observation: Efficiency. For the hard env (As described in  (Carlsson et al.,, 2024)), it seems clear from Figure 12 that though both LATS and LAGEX shows minimally better efficiency than Uniform sampling, LAGEX shows minimum sample compleity in the unknown constraint setup. On the contrary, in easy env the betterment in efficiency for LAGEX is more prominent.
Figure 11: Sample complexity (median±$\pm$±std.) of algorithm for hard env. in Setup 2.
Figure 12: Sample complexity (median±$\pm$±std.) of algorithms for easy env in Setup 2.
Appendix J Technical results and known tools in BAI and pure exploration
@@ I.1 Sample complexity plots for Setup 2
Here, we want to get concentration on the deviation of the pessimistic estimate of the constraint matrix from the actual one quantified by $\|(\tilde{\mathbf{A}}-\mathbf{A})\boldsymbol{\omega}\|_{\infty}$, it becomes very crucial to prove upper bounds on sample complexity of our proposed algorithms. The following lemma ensures the concentration of the constraint matrix.
@@
For the pessimistic estimate $\tilde{\mathbf{A}}$ of $\mathbf{A}$, the following holds
@@
    $|(\tilde{\mathbf{A}}^{i}-\mathbf{A}^{i})\boldsymbol{\omega}|\leq\rho(t,\delta)$ where $\rho(t,\delta)\mathrel{\triangleq}f(t,\delta)\|\boldsymbol{\omega}\|_{\Sigma_{t}^{-1}}$
@@
The first result gives control on the deviations $(\tilde{\mathbf{A}}-\mathbf{A})\boldsymbol{\omega}$ for $\mathbf{A}\in\mathcal{C}_{t}^{\mathbf{A}}(\delta)$. Then $\forall i\in[d]$
$|(\tilde{\mathbf{A}}^{i}-\mathbf{A}^{i})\boldsymbol{\omega}|\leq|(\tilde{\mathbf{A}}_{t}^{i}-\mathbf{A}^{i})\boldsymbol{\omega}|\leq 2\sup_{\mathbf{A}\in\mathcal{C}_{t}^{\mathbf{A}}(\delta)}\|\tilde{\mathbf{A}}_{t}^{i}-\mathbf{A}^{i}\|_{\Sigma_{t}}\|\boldsymbol{\omega}\|_{\Sigma_{t}^{-1}}\leq f(t,\delta)\|\boldsymbol{\omega}\|_{\Sigma_{t}^{-1}}\leq\rho(t,\delta)$
@@
Now we proceed to state an upper bound on the cumulative probability of the bad event i.e the event $|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}|_{\infty}>|(\tilde{\mathbf{A}}_{t}^{i}-\mathbf{A}^{i})\boldsymbol{\omega}|>\rho(t,\delta)$.
@@
The cumulative probability of the bad event till time t<T,
$\sum_{s=1}^{t}\mathbb{P}\left(\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}\|_{\infty}>\rho(t,\delta)\right)\leq\zeta(1)$
@@
We already have stated in the main paper $f(t,\delta)=1+\sqrt{\frac{1}{2}\log\frac{K}{\delta}+\frac{1}{4}\log{\sf det}\Sigma_{t}}\leq 1+\sqrt{\frac{1}{2}\log\frac{K}{\delta}+\frac{d}{4}\log\left(1+\frac{t}{d}\right)}\mathrel{\triangleq}f^{\prime}(t,\delta)$ by Lemma 3 and we define $\rho^{\prime}(t,\delta)\mathrel{\triangleq}f^{\prime}(t,\delta)\|\boldsymbol{\omega}\|_{\Sigma_{t}^{-1}}$. It implies $\mathbb{P}\left(\exists t\in[T]\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}\|_{\infty}>\rho^{\prime}(t,\delta)\right)\leq\mathbb{P}\left(\exists t\in[T]\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}\|_{\infty}>\rho(t,\delta)\right)\leq\delta$. Now if we replace $\log\frac{1}{\delta}$ by u, we can write $\mathbb{P}\left(\exists t\in[T],\forall i\in[d]|\|\tilde{\mathbf{A}}_{t}^{i}-\mathbf{A}^{i}\|_{\Sigma_{t}}>1+\sqrt{\frac{1}{2}\log{K}+\frac{u}{2}+\frac{d}{4}\log\left(1+\frac{t}{d}\right)}\right)\leq\exp(-u)$. We can directly assign $\log t$ as the simplest and natural choice for $u$, since $\sum_{s=1}^{\infty}\frac{1}{t}=\zeta(1)$, $\zeta(.)$ being the Euler-Riemann zeta function. Though this integral is improper, it has a Cauchy principal value as Euler–Mascheroni constant which means $\sum_{s=1}^{\infty}\frac{1}{t}\approx\gamma=0.577$. So we assign $u=\log t$
$\sum_{s=1}^{t}\mathbb{P}\left(\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}\|_{\infty}>\rho(t,\delta)\right)\leq\sum_{s=1}^{t}\frac{1}{t}\leq\sum_{s=1}^{\infty}\frac{1}{t}\leq\zeta(1)\approx 0.577$
@@
For any $\boldsymbol{l}\in\mathcal{L}$ and $\boldsymbol{\omega}\in\Delta_{K}$
$-\boldsymbol{l}^{T}\tilde{\mathbf{A}}\boldsymbol{\omega}\leq\mathcal{D}({\boldsymbol{\mu}},\boldsymbol{\omega},\hat{\mathcal{F}})\psi$
where, $\psi=\frac{||(\tilde{\mathbf{A}}-\mathbf{A})\boldsymbol{\omega}||_{\infty}+\max_{i\in[1,N]}(-\mathbf{A}\boldsymbol{\omega})}{\min_{i\in[1,N]}(-\tilde{\mathbf{A}}\boldsymbol{\omega})}$
@@
$(-\boldsymbol{l}^{T}\tilde{\mathbf{A}}\boldsymbol{\omega})$ $=\boldsymbol{l}^{T}(-\tilde{\mathbf{A}}+\mathbf{A}-\mathbf{A})\boldsymbol{\omega}$
$\leq\|\boldsymbol{l}\|_{1}\|(\mathbf{A}-\tilde{\mathbf{A}})\boldsymbol{\omega}\|_{\infty}+\|\boldsymbol{l}\|_{1}\max_{i\in[1,N]}(-\mathbf{A}^{i}\boldsymbol{\omega})$
$\leq\|\boldsymbol{l}\|_{1}\left((\|\mathbf{A}-\tilde{\mathbf{A}})\boldsymbol{\omega}\|_{\infty}+\max_{i\in[1,N]}(-\mathbf{A}\boldsymbol{\omega})\right)$
$\leq\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\frac{\|(\tilde{\mathbf{A}}-\mathbf{A})\boldsymbol{\omega}\|_{\infty}+\max_{i\in[1,N]}(-\mathbf{A}\boldsymbol{\omega})}{\min_{i\in[1,N]}(-\tilde{\mathbf{A}}\boldsymbol{\omega})}$
@@
Lemma 7.
(Lemma 19 in (Garivier and Kaufmann,, 2016)) There exists two constants B and C (depends on ${\boldsymbol{\mu}}$ and $\epsilon$), such that—
$\sum_{t=h(T)}^{T}\mathbb{P}\left\{\|\hat{{\boldsymbol{\mu}}}_{t}-{\boldsymbol{\mu}}\|_{\infty}>\xi(\epsilon)\right\}\leq BT\exp(-CT^{\frac{1}{8}})$
Lemma 8.
(Garivier and Kaufmann,, 2016, Lemma 7) For all $t\geq 1$ and $\forall a\in[{K}]$, C-Tracking ensures $N_{a,t}\geq\sqrt{t+{K}^{2}}-{K}$ and
@@
Lemma 9.
(Theorem 14 in (Kaufmann and Koolen,, 2021)) Let $\delta>0,\nu$ be independent one-parameter exponential families with mean ${\boldsymbol{\mu}}$ and $S\subset[d]$. Then we have,
@@
$\forall z\in[1,e],\forall x\geq 0,\quad\tilde{h}_{z}(x)$ =$=\begin{cases}\exp\left(\frac{1}{h^{-1}(x)}\right)h^{-1}(x)&\text{ if }x\geq h^{-1}\left(\frac{1}{\ln(z)}\right)\\ z(x-\ln(\ln(z)))&\text{ else }\end{cases}\,.$ (31)
Lemma 10.
(Lemma 17 in (Degenne and Koolen,, 2019)) Under the good event $G_{T}$, there exists a $T_{\epsilon}$ such that for $T$ where $h(T)\geqslant T_{\epsilon}$ C-tracking will satisfy
@@
Lemma 11.
(Theorem 2 in (Degenne et al., 2019b, )) The sample complexity of GE is
$\mathbb{E}[\tau]\leqslant T_{0}(\delta)+\frac{eK}{a}$
where
$T_{0}(\delta)=\max\left\{t\in\mathbb{N}:t\leqslant T(\boldsymbol{{\boldsymbol{\mu}}})c(t,\delta)+C_{\boldsymbol{{\boldsymbol{\mu}}}}\left(R_{t}^{\boldsymbol{{\boldsymbol{\lambda}}}}+R_{t}^{\boldsymbol{w}}+O(\sqrt{t\log t})\right)\right\}$
where $R_{t}^{\boldsymbol{{\boldsymbol{\lambda}}}}$ is the regret of the instance player, $R_{t}^{\boldsymbol{w}}$ the regret of the allocation player and $C_{\boldsymbol{{\boldsymbol{\mu}}}}$ an instance-dependent constant.
J.3 Useful definitions and theorems from literature on continuity of convex functions
Definition 2.
(Definition of Upper Hemicontinuity) We say that a set-valued function $C:\Theta\rightarrow\boldsymbol{\omega}$ is upper hemicontinuous at the point $\theta\in\Theta$ if for any open set $S\subset\boldsymbol{\omega}$ with $C(\theta)\in S$ there exists a neighborhood $U$ around $\theta$, such that $\forall x\in U,C(x)$ is a subset of $S$.
Theorem 8.
(Berge’s maximum theorem, (Berge,, 1963)) Let $X$ and $\Theta$ be topological spaces. Let $f:X\times\Theta\rightarrow\mathbb{R}$ be a continuous function and let $C:\Theta\rightarrow\bar{X}$ be a compact-valued correspondence such that $C(\theta)\neq\varnothing\forall\theta\in\Theta$. If $C$ is continuous at $\theta$ then $f^{*}(\theta)=\sup_{x\in C(\theta)}f(x,\theta)$ is continuous and $C^{*}=\left\{x\in C(\theta):f(x,\theta)=f^{*}(\theta)\right\}$ is upper hemicontinuous.
Theorem 9.
(Heine-Borel theorem, Eduard Heine and Émile Borel) For a subset $S$ in $\mathbb{R}^{n}$, the following two statements are equivalent
@@
(II.1.4). If $H(f)\cap C$ has nonempty interior and no component of $f$ is identically zero, then $\lim_{r\rightarrow\infty}\left(H\left(f^{r}\right)\cap C=H(f)\cap C\right.$
Lemma 12.
(Magureanu et al.,, 2014, Lemma 13) Consider $A\in\left(\mathbb{R}^{+}\right)^{k\times k},c\in\left(\mathbb{R}^{+}\right)^{k}$, and $\mathcal{T}\subset\left(\mathbb{R}^{+}\right)^{k\times k}\times\left(\mathbb{R}^{+}\right)^{k}$. Define $t=(A,c)$. Consider the function $Q$ and the set-valued map $Q^{\star}$

### v2

@@ H.1 Upper Bound on Constraint Violation
In a linear programming problem we say constraint is violated if the chosen allocation fails to satisfy any of the true linear constraints. In other words when the event $A\boldsymbol{\omega}_{t}\geq 0$. We start with the optimization problem relaxed with slack if the constraints were known,
$\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}$
such that, $A\boldsymbol{\pi}+\Gamma\leq 0$
where, $\Gamma$ is the slack. Cumulative violation of constraints can be expressed as $\mathcal{V}_{t}=\sum_{s=1}^{t}\max_{i\in[K]}\left[A^{i}\boldsymbol{\omega}_{t}\right]_{+}$ where $z]_{+}=\max\{z,0\}$. Then, at any time step $t\in[T]$, instantaneous violation is given by, $v_{t}=\max_{i\in[k]}\left[A^{i}\boldsymbol{\omega}_{t}\right]_{+}$. Since, $A$ is feasible, we define the game value as, $\eta=\max_{\boldsymbol{\omega}\in\mathcal{F}}\max_{i\in[K]}A^{i}\boldsymbol{\omega}\leq-\Gamma$ and $\eta=\max_{\boldsymbol{\omega}\in\mathcal{F}}\max_{i\in[K]}A^{i}\boldsymbol{\omega}\geq\min_{\tilde{A}\in\mathcal{C}_{A}^{t}}\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}_{t}}\max_{i\in[K]}\tilde{A}_{t}^{i}\boldsymbol{\omega}_{t}=\tilde{A}_{t}^{i_{t}}\boldsymbol{\omega}_{t}$, holds due to pessimistic estimate of $A$.
Again, we define, $i_{\text{min}}(\boldsymbol{\omega})=\arg\min_{i\in[K]}\tilde{A}^{i}\boldsymbol{\omega}$.
$\max_{i\in[K]}\left[A^{i}\boldsymbol{\omega}_{t}\right]$ $=\max_{i\in[K]}(A^{i}-\tilde{A}_{t}^{i})\boldsymbol{\omega}_{t}+\max_{i\in[K]}\tilde{A}_{t}^{i}\boldsymbol{\omega}_{t}$
[…]
where the last inequality holds because $\max_{i\in[K]}\tilde{A}_{t}^{i}\boldsymbol{\omega}_{t}\leq\max_{i\in[K]}A^{i}\boldsymbol{\omega}_{t}\leq-\Gamma$
@@
$\mathbf{E}\big[\frac{\mathcal{V}_{\tau_{\delta}}}{\tau_{\delta}}\big]=\sum_{t\leq\mathbf{E}[\tau_{\delta}]}s_{t}\leq\sum_{t\leq\mathbf{E}[\tau_{\delta}]}[\rho(t,\delta)-\Gamma]_{+}=$ $\sum_{t\leq\mathbf{E}[\tau_{\delta}]}\{\rho(t,\delta)-\Gamma\}\,\hskip 5.69046pt\text{because }\rho(t,\delta)-\Gamma\hskip 2.84544pt\text{ is positive.}$
@@ H.2 Experimental Results on Constraint Violation
We compare the constraint violation i.e the number of times $A\boldsymbol{\omega}_{t}>0$, where $A$ is the true constraint matrix.
[…]
Observation: LAGEX is the “safest” among the existing baselines.
@@
Therefore, we see that across multiple environments LAGEX is the ”safest” out of all the competing algorithms.
@@ Appendix I Experimental Details
Computation resources. We run the algorithms on a 64-bit 13th Gen Intel® Core™ i7-1370P × 20 processor machine with 32GB ram with a disc capability of 1TB and graphics Mesa Intel® Graphics (RPL-P). The hardware model is HP EliteBook 840 14 inch G10 Notebook PC.
@@
Appendix J Technical Results: New and Known
@@
Here, we want to get concentration on the deviation of the pessimistic estimate of the constraint matrix from the actual one quantified by $\|(\tilde{A}-A)\boldsymbol{\omega}\|_{\infty}$, it becomes very crucial to prove upper bounds on sample complexity of our proposed algorithms. The following lemma ensures the concentration of the constraint matrix.
@@
For the pessimistic estimate $\tilde{A}$ of $A$, the following holds
@@
    $|(\tilde{A}^{i}-A^{i})\boldsymbol{\omega}|\leq\rho(t,\delta)$ where $\rho(t,\delta)\mathrel{\triangleq}f(t,\delta)\|\boldsymbol{\omega}\|_{\Sigma_{t}^{-1}}$
@@
The first result gives control on the deviations $(\tilde{A}-A)\boldsymbol{\omega}$ for $A\in\mathcal{C}_{t}^{A}(\delta)$. Then $\forall i\in[d]$
$|(\tilde{A}^{i}-A^{i})\boldsymbol{\omega}|\leq|(\tilde{A}_{t}^{i}-A^{i})\boldsymbol{\omega}|\leq 2\sup_{A\in\mathcal{C}_{t}^{A}(\delta)}\|\tilde{A}_{t}^{i}-A^{i}\|_{\Sigma_{t}}\|\boldsymbol{\omega}\|_{\Sigma_{t}^{-1}}\leq f(t,\delta)\|\boldsymbol{\omega}\|_{\Sigma_{t}^{-1}}\leq\rho(t,\delta)$
@@
Now we proceed to state an upper bound on the cumulative probability of the bad event i.e the event $|(\tilde{A}_{t}-A)\boldsymbol{\omega}|_{\infty}>|(\tilde{A}_{t}^{i}-A^{i})\boldsymbol{\omega}|>\rho(t,\delta)$.
@@
The cumulative probability of the bad event till time t¡T,
$\sum_{s=1}^{t}\mathbb{P}\left(\|(\tilde{A}_{t}-A)\boldsymbol{\omega}\|_{\infty}>\rho(t,\delta)\right)\leq\zeta(1)$
@@
We already have stated in the main paper
$f(t,\delta)=1+\sqrt{\frac{1}{2}\log\frac{K}{\delta}+\frac{1}{4}\log{\sf det}\Sigma_{t}}\leq 1+\sqrt{\frac{1}{2}\log\frac{K}{\delta}+\frac{d}{4}\log\left(1+\frac{t}{d}\right)}\mathrel{\triangleq}f^{\prime}(t,\delta)$
by Lemma 3 and we define $\rho^{\prime}(t,\delta)\mathrel{\triangleq}f^{\prime}(t,\delta)\|\boldsymbol{\omega}\|_{\Sigma_{t}^{-1}}$. It implies $\mathbb{P}\left(\exists t\in[T]\|(\tilde{A}_{t}-A)\boldsymbol{\omega}\|_{\infty}>\rho^{\prime}(t,\delta)\right)\leq\mathbb{P}\left(\exists t\in[T]\|(\tilde{A}_{t}-A)\boldsymbol{\omega}\|_{\infty}>\rho(t,\delta)\right)\leq\delta$. Now if we replace $\log\frac{1}{\delta}$ by u, we can write
$\mathbb{P}\left(\exists t\in[T],\forall i\in[d]|\|\tilde{A}_{t}^{i}-A^{i}\|_{\Sigma_{t}}>1+\sqrt{\frac{1}{2}\log{K}+\frac{u}{2}+\frac{d}{4}\log\left(1+\frac{t}{d}\right)}\right)\leq\exp(-u).$
We can directly assign $\log t$ as the simplest and natural choice for $u$, since $\sum_{s=1}^{\infty}\frac{1}{t}=\zeta(1)$, $\zeta(.)$ being the Euler-Riemann zeta function. Though this integral is improper, it has a Cauchy principal value as Euler–Mascheroni constant which means $\sum_{s=1}^{\infty}\frac{1}{t}\approx\gamma=0.577$. So we assign $u=\log t$
$\sum_{s=1}^{t}\mathbb{P}\left(\|(\tilde{A}_{t}-A)\boldsymbol{\omega}\|_{\infty}>\rho(t,\delta)\right)\leq\sum_{s=1}^{t}\frac{1}{t}\leq\sum_{s=1}^{\infty}\frac{1}{t}\leq\zeta(1)\approx 0.577$
@@
For any $\boldsymbol{l}\in\mathcal{L}$ and $\boldsymbol{\omega}\in\Delta_{K}$, $\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\leq\mathcal{D}({\boldsymbol{\mu}},\boldsymbol{\omega},\hat{\mathcal{F}})\psi$
[…]
@@
$(-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega})=\boldsymbol{l}^{\top}(-\tilde{A}+A-A)\boldsymbol{\omega}$ $\leq\|\boldsymbol{l}\|_{1}\|(A-\tilde{A})\boldsymbol{\omega}\|_{\infty}+\|\boldsymbol{l}\|_{1}\max_{i\in[1,N]}(-A^{i}\boldsymbol{\omega})$
$\leq\|\boldsymbol{l}\|_{1}\left((\|A-\tilde{A})\boldsymbol{\omega}\|_{\infty}+\max_{i\in[1,N]}(-A\boldsymbol{\omega})\right)$
$\leq\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F})\frac{\|(\tilde{A}-A)\boldsymbol{\omega}\|_{\infty}+\max_{i\in[1,N]}(-A\boldsymbol{\omega})}{\min_{i\in[1,N]}(-\tilde{A}\boldsymbol{\omega})}$
@@
Lemma 7 (Lemma 19 in (Garivier and Kaufmann,, 2016)).
There exists two constants B and C (depends on $\boldsymbol{\mu}$} and $\epsilon$), such that—
$\sum_{t=h(T)}^{\top}\mathbb{P}\left\{\|\hat{{\boldsymbol{\mu}}}_{t}-{\boldsymbol{\mu}}\|_{\infty}>\xi(\epsilon)\right\}\leq BT\exp(-CT^{\frac{1}{8}})$
Lemma 8 (Lemma 7, Garivier and Kaufmann, (2016)).
For all $t\geq 1$ and $\forall a\in[{K}]$, C-Tracking ensures $N_{a,t}\geq\sqrt{t+{K}^{2}}-{K}$ and
@@
Lemma 9 (Theorem 14 in (Kaufmann and Koolen,, 2021)).
Let $\delta>0,\nu$ be independent one-parameter exponential families with mean $\boldsymbol{\mu}$} and $S\subset[d]$. Then we have,
@@
[…]
Lemma 10 (Lemma 17 in (Degenne and Koolen,, 2019)).
Under the good event $G_{T}$, there exists a $T_{\epsilon}$ such that for $T$ where $h(T)\geqslant T_{\epsilon}$ C-tracking will satisfy
@@
Lemma 11 (Theorem 2 in (Degenne et al., 2019b, )).
The sample complexity of GE is $\mathbb{E}[\tau]\leqslant T_{0}(\delta)+\frac{eK}{a}$ where $T_{0}(\delta)=\max\left\{t\in\mathbb{N}:t\leqslant T(\boldsymbol{{\boldsymbol{\mu}}})c(t,\delta)+C_{\boldsymbol{{\boldsymbol{\mu}}}}\left(R_{t}^{\boldsymbol{{\boldsymbol{\lambda}}}}+R_{t}^{\boldsymbol{w}}+O(\sqrt{t\log t})\right)\right\}$ where $R_{t}^{\boldsymbol{{\boldsymbol{\lambda}}}}$ is the regret of the instance player, $R_{t}^{\boldsymbol{w}}$ the regret of the allocation player and $C_{\boldsymbol{{\boldsymbol{\mu}}}}$ an instance-dependent constant.
J.3 Useful Results for Continuity of Convex Functions
Definition 2 (Definition of Upper Hemicontinuity).
We say that a set-valued function $C:\Theta\rightarrow\boldsymbol{\omega}$ is upper hemicontinuous at the point $\theta\in\Theta$ if for any open set $S\subset\boldsymbol{\omega}$ with $C(\theta)\in S$ there exists a neighborhood $U$ around $\theta$, such that $\forall x\in U,C(x)$ is a subset of $S$.
Theorem 8 (Berge’s maximum theorem, (Berge,, 1963)).
Let $X$ and $\Theta$ be topological spaces. Let $f:X\times\Theta\rightarrow\mathbb{R}$ be a continuous function and let $C:\Theta\rightarrow\bar{X}$ be a compact-valued correspondence such that $C(\theta)\neq\varnothing\forall\theta\in\Theta$. If $C$ is continuous at $\theta$ then $f^{*}(\theta)=\sup_{x\in C(\theta)}f(x,\theta)$ is continuous and $C^{*}=\left\{x\in C(\theta):f(x,\theta)=f^{*}(\theta)\right\}$ is upper hemicontinuous.
Theorem 9 (Heine-Borel theorem, due to Eduard Heine and Émile Borel).
For a subset $S$ in $\mathbb{R}^{n}$, the following two statements are equivalent
@@
(II.1.4). If $H(f)\cap C$ has nonempty interior and no component of $f$ is identically zero, then $\lim_{r\rightarrow\infty}\left(H\left(f^{r}\right)\cap C\right)=H(f)\cap C.$
Lemma 12 (Lemma 13, Magureanu et al., (2014)).
Consider $A\in\left(\mathbb{R}^{+}\right)^{k\times k},c\in\left(\mathbb{R}^{+}\right)^{k}$, and $\mathcal{T}\subset\left(\mathbb{R}^{+}\right)^{k\times k}\times\left(\mathbb{R}^{+}\right)^{k}$. Define $t=(A,c)$. Consider the function $Q$ and the set-valued map $Q^{\star}$